sun. (3 (3pts) MY S +2 9 25. +
Answer #1

Given that 2구 2 + 뜻 + + t. 16 + it is GP geries t tom 으 2 = 9 (common ratio = 2 =21 Sn 이나에) I-r sn=2다1M 1- 2. Sn = 21) 3 아 Sn